# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v2

@@ 4.2. Global Arch: Towards Behavior-Aware Interaction Learning
By performing interaction learning on the concatenation of $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$, Eq. (7) incorporates three different interactions: global-global, global-sequence, and sequence-sequence interactions. The global-global interactions capture explicit user interests by computing the relevance between the user profile and the target item content, while the global-sequence interactions capture implicit user interests by computing the relevance between the target item and the user’s current need within the interaction sequence. Additionally, by computing the interactions among sequence features, the most representative information is promoted with high interaction scores, while inactive items with low scores are filtered out. By stacking multiple layers, the Global Arch captures rich behavior-aware interactions at different orders.
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
In addition to the explicit static user preference in global features, the implicit dynamic user interests behind the behavior sequences provides complementary information (Zhou et al., 2019; Sun et al., 2019). However, given that user behavior sequence can be highly random and noisy, solely relying on sequence information is ineffective. It is important to incorporate global context into sequence modeling. For example, a user may randomly browse items on online shopping platforms, but given the global information that the user is an electronic fan, electronic items, e.g., smartphone and laptops, in the browsing history are identified as key information that requires extra attention.
To learn context-aware sequence embeddings, a Sequence Arch is designed based on two key ideas: Personalized FFN (PFFN) and Multi-Head Attention (MHA). To enable interactions between global and sequence information, we first employ PFFN to transform sequence embeddings given global summarization as query. Given global summarization $\mathbf{X}_{\text{sum}}^{(l)}$ and sequence embedding $\mathbf{S}^{(l)}$ at the $l$-th layer, the PFFN operation is defined as:
@@
where $f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)$ is an MLP that aims to learn the linear projection on sequence based on global summarization.
Besides, to model the relationship among events in a sequence, MHA is applied to enable the model to attend to different parts of the sequence, capturing long-range dependencies and contextual information. To incorporate global context, before feeding into the first InterFormer layer, the global summarization $\mathbf{X}^{(1)}_{\text{sum}}$ is prepended to the sequence as the CLS token, i.e., $\mathbf{S}^{(1)}=\left[\mathbf{X}^{(1)}_{\text{sum}}\|\mathbf{S}^{(1)}\right]$, hence the following MHA on the CLS token can aggregate sequence information using global information as the query (Vaswani et al., 2017; Devlin et al., 2018). This is similar to the early fusion idea described in Transact  (Xia et al., 2023) but we primarily consider ’append’ instead of ’concat’ as the fusion method. Additionally, rotary position embeddings (Su et al., 2024) are applied on tokens so that the positional information in sequences can be effectively leveraged. In general, the Sequence Arch can be written as follow:
@@
Since the output $\mathbf{S}^{(l+1)}$ is of the same shape as the input $\mathbf{S}^{(l)}$, aggressive sequence summarization can be avoided and layers can be easily stacked.
Through layer stacking, on one hand, the sequence embeddings is aware of global interactions at different orders via PFFN; on the other hand, MHA at different layers focuses on different parts of the sequence, capturing multi-scale sequence information. Therefore, the model can learn a rich context-aware encoding of the sequential data capturing both local and global patterns within the sequence.
4.4. Bridging Arch: Towards Effective Information Selection and Summarization
Though selective information aggregation is achieved in both the global and sequence arch, as the dimensionalities of input and output features are retained till the final layer, it is infeasible to directly exchange such information due to (1) ineffectiveness given the noisy information and (2) inefficiency given the high-dimensionality. To solve this dilemma, we propose a bridging arch to select and summarize information before exchanging between different data modes.
To start with, given the large scale of global embeddings, it is important to selectively summarize them to guide sequence learning. To this end, we highlight the most useful information through a personalized gated selection mechanism as follows:
@@ 4.4. Bridging Arch: Towards Effective Information Selection and Summarization
where $\text{where }\text{Gating}(\mathbf{X})=\mathbf{X}\odot\text{MLP}(\mathbf{X}).$
where MLP: $\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times n_{\text{sum}}}$ with $n_{\text{sum}}\ll n$, and self-gating (Chai et al., 2020) provides sparse masking on the embeddings such that relevant information is retained while irrelevant noises are filtered out, providing high-quality context for efficient sequence learning.
For sequence information, three types of summarization are neatly designed, including CLS tokens, PMA tokens and recent interacted tokens. The CLS tokens $\mathbf{S}_{\text{CLS}}$ learned by MHA are selected as context-aware sequence summarization. However, the quality of CLS tokens largely depend on the learned global context. To compensate for the heavy reliance on global context and enable more flexibility, the PMA tokens $\mathbf{S}_{\text{PMA}}$ (Lee et al., 2019), which are essentially sequence summarization based on learnable queries, are employed. Besides, the $K$ most recent interacted tokens ($\mathbf{S}_{\text{recent}}$) have been proven to be effective in capturing the user’s recent interests (Borisyuk et al., 2024; Xia et al., 2023). The combination of the above information is further gated by a self-gating layer, serving as the behavior summarization for global interaction:
@@
In general, the benefits of the bridging arch are two-fold. On the one hand, by separating information summarization from global and sequence arch, information can be retained in both arches, avoiding aggressive information aggregation. On the other hand, effective information exchange can be achieved as high-dimensional global/sequence features are selected and summarized into low-dimensional embeddings. Therefore, the Bridging Arch plays a pivotal role in enabling the model to capture complex interactions between global behaviors and sequential patterns, leading to more comprehensive representations of the input data.
@@ 5. Experiment
We evaluate the proposed InterFormer to answer the following research questions:
  • Q1: How effective is the proposed InterFormer (Section 5.2)?
  • Q2: How scalable is the proposed InterFormer (Section 5.3)?
  • Q3: To what extent does InterFormer address the insufficient inter-mode interaction and aggressive information aggregation (Section 5.4)?
  • Q4: To what extent does InterFormer benefit from different modules (Section 5.5)?
@@ 5.1. Experiment Setup
Datasets. We adopt three benchmark benchmark datasets, including AmazonElectronics (He and McAuley, 2016b), TaobaoAd (Tianchi, [n. d.]), KuaiVideo (Li et al., 2019), and a large-scale internal dataset for evaluation. Dataset statistics are summarized in Table 2 with more details in Appendix B.1.
@@
Dataset #Samples #Feat. (Seq/Global) Seq Length
@@
Baseline Methods. We compare the proposed InterFormer with 11 state-of-the-art models, including (1) non-sequential methods: FM (Rendle, 2010), xDeepFM (Lian et al., 2018), AutoInt+ (Song et al., 2019), DCNv2 (Wang et al., 2021b), FmFM (Sun et al., 2021), DOT product, DHEN (Zhang et al., 2022), Wukong (Zhang et al., 2024), and (2) sequential methods: DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), BST (Chen et al., 2019), DMIN (Xiao et al., 2020), DMR (Lyu et al., 2020), TransAct (Xia et al., 2023). We adopt DHEN as the global arch for InterFormer in the experiments. Detailed model configurations are provided in Appendix B.2.
@@
5.2. Effectiveness Results
@@
The experiment results are shown in Table 3. We first observe that sequential methods consistently outperforms non-sequential methods on all datasets. For non-sequential methods, sequence information is naively aggregated in early stages and further processed together with global information. While in sequential methods, neatly designed sequence processing modules, e.g., RNN (Zhou et al., 2019; Sun et al., 2019) and attention mechanism (Zhou et al., 2018; Chen et al., 2019; Xiao et al., 2017; Lyu et al., 2020), are employed, so that sequential information can be processed in aware of global context. The universal outperformance of sequential methods validates that different data modes should be processed differently. Besides, results show that aggressive sequence summarization in the early stages, without considering global context, will impair model performance.
5.3. Scalability Results
To validate the scalability of InterFormer, we carry out experiments on a large-scale internal dataset from Meta, containing 70B samples in total, hundreds of non-sequence features and 10 sequences of length 200 to 1,000. In general, a 3-layer InterFormer achieves a 0.15% NE gain compared to the internal SOTA model with similar FLOPs and 24% Queries Per Second (QPS) gain. Together with feature scaling, the improvement on NE can be further enlarged with a 10% of Model FLOPs Utilization (MFU) on 512 GPUs attained. InterFormer shows great generalizability on a wide range of models showing promising ROI.
On Sequence Feature Scaling. We evaluate how the performance of InterFormer changes when the sequence feature scale increases in internal dataset. In addition to six sequences of length 100, we include two additional long sequences of length 1000, and we observe a 0.14% improvement in NE. As shown in Figure 2(a), InterFormer exhibits better scalability compared to the strong internal baseline (Internal) that leverages cross-attention to capture sequence and non-sequence information, as NE curve of InterFormer continues to decrease when more training samples are involved, outperforming Internal by 0.06% in NE. Besides, we also tried to merge six sequences to generate one long sequence of length 600, and we observe promising efficiency improvements in QPS (+20%) and MFU (+17%) with a 0.02% NE tradeoff. From the modeling perspective, the results validate that InterFormer is able to enlarge the NE gain brought by the sequence feature scaling.
(a) Feature scaling
(b) Model scaling
On Model Scaling. In addition to feature scaling, we evaluate how InterFormer performs when the model scale increases. As Figure 2(b) shows, scaling InterFormer from 1 to 4 layers achieves consistent NE gains, exhibiting good scaling properties. Specifically, compared to a single layer InterFormer, a two-layer InterFormer achieves a significant 0.13% NE gain, with an additional 0.05% and 0.04% NE gain achieved by stacking 3 and 4 layers, respectively.
Model-System Co-Design. Model architecture design plays a crucial role in training efficiency with its implication on GPU FLOPs utilization and inter-GPU communication. We highlight two optimizations based on InterFormer architecture that in total boost training efficiency by more than 30%, including communication overhead reduction and computation efficiency.
For communication overhead reduction, the global arch DHEN with heavy parameters and relatively light computation tends to be FSDP (Zhao et al., 2023) communication-bound in distributed training, while the sequence arch Transformer with much higher computation (FLOPs) to parameter ratio is normally computation-bound. To alleviate such inefficiency, our parallel design of global and sequence arch allows the exposed communication from DHEN modules to overlap with sequence computation effectively, resulting in a 20% QPS improvement compared to the performing two modules sequentially.
5.4. Analysis of the Model
On the Interleaving Learning Style.
@@ 5.4. Analysis of the Model
We first analyze how the interleaving learning style benefits heterogeneous interaction learning. To validate the universality of InterFormer, we consider three backbone global architectures, including dot product (DOT), DCNv2, and DHEN, and five different scenarios, including
  • sole where only global arch is adopted, while sequence information is naively aggregated in early stage.
  • sep where inter-mode interaction is disabled and sequence and global arch are learnt separately.
  • s2g where only sequence-to-global information flow is enabled, while the reverse direction is disabled.
  • g2s where only global-to-sequence information flow is enabled, while the reverse direction is disabled.
@@
As shown in Figure 3, int consistently outperforms other scenarios regardless of the backbone global arch, with an up to 1.46% outperformance in AUC. This validates the universal benefits brought by our proposed interleaving learning style.
@@
On the Selective Information Aggregation. We evaluate the effect of selective information aggregation by comparing with three aggressive aggregation variants, where sequence information is first compressed by (1) average pooling, (2) MLP, and (3) MHA, and further combined with global features by the interaction module DHEN. The experimental results are shown in Figure 4.
@@
On Sequence Modeling. We also visualize the learned attention map in Figure 5 to understand sequence modeling.
@@
First, we observe that the attention map exhibits a cluster structure, i.e., surrounding tokens are likely to interact, which can be regarded as a weighted pooling within selected surrounding tokens. Such selective pooling process alleviates random noises in the behavior history to extract better temporal user interests.
Secondly, we find that MHAs at different layers focus on different tokens at different scales. For example, the attention map is more uniform in the first layer, indicating pooling in a broad scale for long-term interests. In contrast, the third layer exhibits stronger interactions within a smaller cluster, resulting in fine-grained aggregation for short-term interests. Furthermore, the second layer focuses on specific tokens, e.g., the recent tokens (tokens in right columns) and the 24-th token, for item-specific information.
5.5. Ablation Study
@@
To better understand how different model designs contribute to the final performance, we ablate different exchanged information, including PMA tokens and recent interacted tokens, as well as different modules like the gating module, PFFN and MHA.
As the results shown in Figure 6, all the exchanged information and modules contribute to the model performance. Specifically, PMA tokens provide effective sequence summarization given global context and are of great importance, ablating which decreases AUC by up to 0.004. The gating modules help selective summarize information from different data modes contribute up to 0.003 AUC improvement. The sequence modeling modules (PFFN and MHA) also improves the performance to some extent.

### v3

@@ 4.2. Interaction Arch: Towards Behavior-Aware Interaction Learning
By performing interaction learning on the concatenation of $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$, Eq. (7) incorporates non-sequence and sequence features. The interactions among non-sequence features capture explicit user interests by computing the relevance between user profile and target item content, while the interactions between non-sequence and sequence capture implicit user interests by computing the relevance between target item and user’s current need within the interaction sequence. Additionally, by computing the interactions among sequence features, the most representative information is promoted with high interaction scores, while inactive items with low scores are filtered out. By stacking multiple layers, the Interaction Arch captures rich behavior-aware interactions at different orders.
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
In addition to the explicit static user preference in non-sequence features, the implicit dynamic user interests behind behavior sequences provide complementary information (Zhou et al., 2019; Sun et al., 2019). However, given that behavior sequences are highly random and noisy, solely relying on sequential information is ineffective. It is important to incorporate non-sequential context into sequence modeling. For example, a user may randomly browse items on shopping platforms, but given the static information that the user is an electronic fan, electronic items, e.g., smartphones and laptops, in the browsing history provide key information demanding extra attention.
To learn context-aware sequence embeddings, a Sequence Arch is designed based on two key ideas: Personalized FFN (PFFN) and Multi-Head Attention (MHA). To enable interactions between non-sequential and sequential information, PFFN is employed to transform sequence embeddings given non-sequence summarization as query. Given non-sequence summarization $\mathbf{X}_{\text{sum}}^{(l)}$ and sequence embedding $\mathbf{S}^{(l)}$ at the $l$-th layer, the PFFN operation is defined as:
@@
where $f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)$ is an MLP that aims to learn the linear projection on sequence based on non-sequence summarization.
Besides, to model the relationship among events in a sequence, MHA is applied to enable the model to attend to different parts of the sequence, capturing long-range dependencies and contextual information. To incorporate non-sequential context, before feeding into the first InterFormer layer, the non-sequence summarization $\mathbf{X}^{(1)}_{\text{sum}}$ is prepended to the sequence as the CLS token, i.e., $\mathbf{S}^{(1)}=\left[\mathbf{X}^{(1)}_{\text{sum}}\|\mathbf{S}^{(1)}\right]$, hence the following MHA on the CLS token can aggregate sequential information using non-sequential information as the query (Vaswani et al., 2017; Devlin et al., 2018). This is similar to the early fusion idea described in Transact  (Xia et al., 2023) but we primarily consider ’append’ instead of ’concat’ as the fusion method. Additionally, rotary position embeddings (Su et al., 2024) are applied on tokens so that the positional information in sequences can be effectively leveraged. In general, the Sequence Arch can be written as follow:
@@
As the output $\mathbf{S}^{(l+1)}$ is of the same shape as the input $\mathbf{S}^{(l)}$, aggressive aggregation can be avoided and layers can be easily stacked.
Through layer stacking, for one thing, the sequence embeddings is aware of non-sequence interactions at different orders via PFFN; for another, MHA at different layers focuses on different parts of the sequence, capturing multi-scale sequential information. Therefore, the model can learn a rich context-aware encoding of the sequential data capturing both local and global patterns within the sequence.
4.4. Cross Arch: Towards Effective Information Selection and Summarization
Though selective information aggregation is achieved in both the Interaction and Sequence Arch, as the dimensionalities of input and output features are retained till the final layer, it is infeasible to directly exchange such information due to (1) ineffectiveness given the noisy information and (2) inefficiency given the high-dimensionality. To solve this dilemma, we propose a cross Arch to select and summarize information before exchanging information.
To start with, given the large scale of non-sequence embeddings, it is important to selectively summarize them to guide sequence learning. To this end, we highlight the most useful information through a personalized gated selection mechanism as follows:
@@ 4.4. Cross Arch: Towards Effective Information Selection and Summarization
where $\text{where }\text{Gating}(\mathbf{X})=\sigma\left(\mathbf{X}\odot\text{MLP}(\mathbf{X})\right).$
where MLP: $\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times n_{\text{sum}}}$ with $n_{\text{sum}}\ll n$, and $\sigma$ is the activation function, e.g., sigmoid, tanh or even identity functions. Self-gating (Chai et al., 2020) provides sparse masking on the embeddings such that relevant information is retained while irrelevant noises are filtered out, providing high-quality context for sequence learning.
For sequential information, three types of summarization are neatly designed, including CLS tokens, PMA tokens and recent interacted tokens. The CLS tokens $\mathbf{S}_{\text{CLS}}$ learned by MHA are selected as context-aware sequence summarization. However, the quality of CLS tokens largely depend on the learned non-sequential context. To compensate such heavy reliance and enable more flexibility, the PMA tokens $\mathbf{S}_{\text{PMA}}$ (Lee et al., 2019), which are essentially sequence summarization based on learnable queries, are employed. Besides, the $K$ most recent interacted tokens ($\mathbf{S}_{\text{recent}}$) have been proven to be effective in capturing the user’s recent interests (Borisyuk et al., 2024; Xia et al., 2023). The combination of the above information is further gated by a self-gating layer, serving as the behavior summarization for non-sequence interaction:
@@
In general, the benefits of the cross Arch are two-fold. On the one hand, by separating information summarization from Interaction and Sequence Arch, information can be retained in both arches, avoiding aggressive information aggregation. On the other hand, effective information exchange can be achieved as high-dimensional non-sequence/sequence features are selected and summarized into low-dimensional embeddings. Therefore, the Cross Arch plays a pivotal role in enabling the model to capture complex interactions between non-sequence behaviors and sequential patterns, leading to more comprehensive representations of the input data.
@@ 5. Experiment
In this section, we carry out extensive experiments to evaluate the proposed InterFormer. We first introduce the experiment setup in Section 5.1. Afterwards, we evaluate InterFormer on public benchmark datasets are carried out in Section 5.2. The proposed InterFormer is deployed at Meta, and the post-launch results on internal industrial-scale dataset are provided in Section 5.3.
@@ 5.1. Experiment Setup
Datasets. We adopt three benchmark datasets, including AmazonElectronics (He and McAuley, 2016b), TaobaoAd (Tianchi, [n. d.]), KuaiVideo (Li et al., 2019), and a large-scale internal dataset for evaluation. Dataset statistics are summarized in Table 2 with more details in Appendix B.1.
@@
Dataset #Samples #Feat. (Seq/Non-Seq) Seq Length
@@
Baseline Methods. We compare the proposed InterFormer with 11 state-of-the-art models, including (1) non-sequential methods: FM (Rendle, 2010), xDeepFM (Lian et al., 2018), AutoInt+ (Song et al., 2019), DCNv2 (Wang et al., 2021b), FmFM (Sun et al., 2021), DOT product, DHEN (Zhang et al., 2022), Wukong (Zhang et al., 2024), and (2) sequential methods: DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), BST (Chen et al., 2019), DMIN (Xiao et al., 2020), DMR (Lyu et al., 2020), TransAct (Xia et al., 2023). We adopt DHEN as the Interaction Arch in the experiments. Detailed model configurations are provided in Appendix B.2.
@@
5.2. Evaluation on Benchmark Datasets
5.2.1. Results.
The experiment results are shown in Table 3. We first observe that sequential methods consistently outperforms non-sequential methods on all datasets. For non-sequential methods, sequence information is naively aggregated in early stages and further processed together with non-sequence information. While in sequential methods, neatly designed sequence processing modules, e.g., RNN (Zhou et al., 2019; Sun et al., 2019) and attention mechanism (Zhou et al., 2018; Chen et al., 2019; Xiao et al., 2017; Lyu et al., 2020), are employed, so that sequential information can be processed in aware of non-sequence context. The universal outperformance of sequential methods validates that different data modes should be processed differently. Besides, results show that aggressive sequence summarization in the early stages, without considering non-sequence context, will impair model performance.
@@ 5.2. Evaluation on Benchmark Datasets
5.2.2. Analysis of the Model.
@@
First (Interleaving Learning Style), we analyze how the interleaving learning style benefits heterogeneous interaction learning. To validate the universality of InterFormer, we consider three backbone Interaction Arch, including dot product (DOT), DCNv2, and DHEN, and five different scenarios, including
  • sole where only Interaction Arch is adopted, while sequence information is naively aggregated in early stage.
  • sep where inter-mode interaction is disabled and Sequence and Interaction Arch are learnt separately.
  • s2n where only sequence-to-non-sequence information flow is enabled, while the reverse direction is disabled.
  • n2s where only non-sequence-to-sequence information flow is enabled, while the reverse direction is disabled.
@@
Second (Selective Information Aggregation), we evaluate the effect of selective information aggregation by comparing with three aggressive aggregation variants, where sequence information is first compressed by (1) average pooling, (2) MLP, and (3) MHA, and further combined with non-sequence features by the interaction module DHEN. The experimental results are shown in Figure 3.
@@
Third (Sequence Modeling), we visualize the learned attention map in Figure 4 to understand sequence modeling. We observe that the attention map exhibits a cluster structure, i.e., surrounding tokens are likely to interact, which can be regarded as a weighted pooling within selected surrounding tokens. Such selective pooling process alleviates random noises in the behavior history to extract better temporal user interests. Besides, MHAs at different layers focus on different tokens at different scales. For example, the attention map is quite uniform in the first layer, indicating pooling in a broad scale for long-term interests. In contrast, the third layer exhibits stronger interactions within smaller clusters, resulting in fine-grained aggregation for short-term interests. Furthermore, the second layer focuses on specific tokens, e.g., recent tokens (tokens in right columns) and the 24-th token, for item-specific information.
@@
5.2.3. Ablation Study
@@
To better understand how different model designs contribute to the final performance, we ablate different exchanged information, including PMA tokens and recent interacted tokens, as well as different modules like the gating module, PFFN and MHA. As the results shown in Figure 5, all the exchanged information and modules contribute to the model performance. Specifically, PMA tokens provide effective sequence summarization given non-sequence context and are of great importance, ablating which decreases AUC by up to 0.004. The gating modules help selective summarize information from different data modes contribute up to 0.003 AUC improvement. The sequence modeling modules (PFFN and MHA) also improves the performance to some extent.
5.3. Evaluation on Industrial Datasets
To evaluate InterFormer’s performance at industrial-scale data, we carry out experiments on a large-scale internal dataset from Meta, containing 70B samples in total, hundreds of non-sequence features and 10 sequences of length 200 to 1,000.
5.3.1. Results.
In general, a 3-layer InterFormer achieves a 0.15% NE gain compared to the internal SOTA model with similar FLOPs and 24% Queries Per Second (QPS) gain. Together with feature scaling, the improvement on NE can be further enlarged with a 10% of Model FLOPs Utilization (MFU) on 512 GPUs attained. InterFormer shows great generalizability on a wide range of models showing promising ROI.
Besides (Sequence Feature Scaling), we evaluate how the performance of InterFormer changes when the sequence feature scale increases in internal dataset. In addition to six sequences of length 100, we include two additional long sequences of length 1000, and we observe a 0.14% improvement in NE. As shown in Figure 6(a), InterFormer exhibits better scalability compared to the strong internal baseline (Internal) that leverages cross-attention to capture sequence and non-sequence information, as NE curve of InterFormer continues to decrease when more training samples are involved, outperforming Internal by 0.06% in NE. Besides, we also tried to merge six sequences to generate one long sequence of length 600, and we observe promising efficiency improvements in QPS (+20%) and MFU (+17%) with a 0.02% NE tradeoff. From the modeling perspective, the results validate that InterFormer is able to enlarge the NE gain brought by the sequence feature scaling.
(a) Feature scaling
(b) Model scaling
In addition (Model Scaling), we evaluate how InterFormer performs when the model scale increases. As Figure 6(b) shows, scaling InterFormer from 1 to 4 layers achieves consistent NE gains, exhibiting good scaling properties. Specifically, compared to a single layer InterFormer, a two-layer InterFormer achieves a significant 0.13% NE gain, with an additional 0.05% and 0.04% NE gain achieved by stacking 3 and 4 layers, respectively.
5.3.2. Model-System Co-Design.
Model architecture design plays a crucial role in training efficiency with its implication on GPU FLOPs utilization and inter-GPU communication. We highlight two optimizations based on InterFormer architecture that in total boost training efficiency by more than 30%, including communication overhead reduction and computation efficiency.
For communication overhead reduction, the Interaction Arch DHEN with heavy parameters and relatively light computation tends to be FSDP (Zhao et al., 2023) communication-bound in distributed training, while the Sequence Arch Transformer with much higher computation (FLOPs) to parameter ratio is normally computation-bound. To alleviate such inefficiency, our parallel design of Interaction and Sequence Arch allows the exposed communication from DHEN modules to overlap with sequence computation effectively, resulting in a 20% QPS improvement compared to the performing two modules sequentially.
5.3.3. Online Impact
Interformer has been deployed in key Ads models, including the largest ones through (Luo et al., [n. d.]), showing superior ROI in scaling compared to other state-of-the-art architectures. Pilot launches in 2024 have yielded a 0.6% improvement in topline metrics. We expect great impact to be delivered in subsequent deployment.